how much fraction of area of one face is covered by carbon  atoms in diamond

Grade:12

1 Answers

Gaurav
8 years ago
Diamond has FCC structure
For FCC crystal



Let ‘r’ be the radius of sphere and ‘a’ be the edge length of the cube

As there are 4 sphere in fcc unit cell

∴Volume of four spheres = 4 (4/3πr3)

In fcc, the corner spheres are in touch with the face centred sphere. Therefore, face diagonal AD is equal to four times the radius of sphere

AC= 4r
But from the right angled triangle ACDAC = √AD2+ DC2=√a2+a2=√2a

4r = √2a

or a = 4/√2 r
Packng Fraction = 75%

∴volume of cube = (2/√2 r)3